# The Gauss map of a projective foliation

### v2

@@ 3. Since $deg(\mathcal{F}_{\omega})=n-1:=d$ we get
Foliations of degree two on $\mathbb{P}^{3}$. The irreducible components of the space of foliations of degree two on $\mathbb{P}^{n}$, $n\geq 3$, were classified in [9]. There it is proved that this space has exactly six irreducible components. Using theorem 3 it is not difficult to see that there are two of these components for which the generic element has birational Gauss map: the exceptional component of example 7 and the logarithmic component $\mathcal{L}(1,1,1,1;3)$ of example 3.
@@
This calculation can also be found in 2.5 of [19]. We observe that $\mathcal{N}(2,3)=\{0,1,2,3,4\}$.
@@
Another example is $\mathcal{L}(1,3;3)$ whose Gauss map is birational, is defined in homogeneous coordinates by the form $\omega=3\,f\,dz-z\,df$, where $f(x,y,z,t)=xz^{2}+yzt+y^{3}$.
In this case, we have $Sing(\mathcal{F}_{\omega})=(y=z=0):=V$ and $\mu(\mathcal{F}_{\omega},V)=6$, so that its Gauss map, $\mathcal{G}_{\omega}$, is birational by theorem 3. We have:
@@
The foliation associated with this map has the rational first integral:
@@
then this is also a self-dual foliation.
Example 10.
Foliations of degree two on $\mathbb{P}^{n}$, $n\geq 4$. Here we will see that the Gauss map of the generic foliation of degree two on $\mathbb{P}^{n}$ is not birational if $n\geq 4$. According to [9] the space of foliations of degree two on $\mathbb{P}^{n}$, $n\geq 3$, has six irreducible components:
  • (a).
    $PB(2\,;n)$; the component containing pull-back of foliations of degree two on $\mathbb{P}^{2}$ by a linear map $L\colon\mathbb{P}^{n}\dasharrow\mathbb{P}^{2}$.
  • (b).
  • (c).
    $\mathcal{L}(1,1,1,1\,;n)$; component of foliations defined by logarithmic forms of the type $\sum_{j=1}^{4}\lambda_{j}\frac{d\ell_{j}}{\ell_{j}}$, where $deg(\ell_{j})=1$, $1\leq j\leq 4$.
  • (d).
    $\mathcal{L}(1,1,2\,;n)$; component of foliations defined by logarithmic forms of the type $\lambda_{1}\frac{d\ell_{1}}{\ell_{1}}+\lambda_{2}\frac{d\ell_{2}}{\ell_{2}}+\lambda_{3}\frac{dq}{q}$, where $deg(\ell_{1})=deg(\ell_{2})=1$ and $deg(q)=2$.
  • (e).
    $\mathcal{L}(1,3\,;n)$; component of foliations with a first integral of the form $f/\ell^{3}$, where $deg(\ell)=1$ and $deg(f)=3$.
  • (f).
    $\mathcal{L}(2,2\,;n)$; component of foliations with a first integral of the form $q_{1}/q_{2}$, where $deg(q_{1})=deg(q_{2})=2$.
Examples (a), (b), and (c), cannot be dominant on $\mathbb{P}^{n}$, $n\geq 4$. For instance, an exceptional foliation on $\mathbb{P}^{n}$, $n\geq 4$, is a linear pull-back of the exceptional foliation $\mathcal{E}$ on $\mathbb{P}^{3}$ (cf. [9]), and so its Gauss map cannot be dominant.
Generic foliations in components (d), (e), and (f), have dominant Gauss maps, but not birational. For instance, in case (e) we can assume that $f$ is a generic cubic polynomial. In homogeneous coordinates $(s,t,x,y,z)$ we can suppose that $\ell=s$, so that in the affine coordinate system $s=1$ the foliation $\mathcal{F}\in\mathcal{R}(1,3)$ has the first integral $f(1,t,x,y,z)$. The hyperplane $(t=c)$ is tangent to $\mathcal{F}$ at the points for which $f_{x}(1,c,x,y,z)=f_{y}(1,c,x,y,z)=f_{z}(1,c,x,y,z)=0$. If $f$ and $c$ are generic this set contains $8$ points and so the Gauss map cannot be birational.
$f(s,t,x,y,z)=\frac{x^{2}+y^{2}+z^{2}}{st}:=\frac{q}{st}$
which is defined in homogeneous coordinates by the form
The Gauss map in the affine coordinate system $t=1$ is
and has the inverse
$\mathcal{G}_{\mathcal{F}}^{-1}(a,b,c,e)=\left(-\frac{4a}{q(a,b,c)},-\frac{4b}{q(a,b,c)},-\frac{4c}{q(a,b,c)},\frac{1}{e}\right)$
Next, we will consider foliations defined by closed forms. Among them we have the logarithmic foliations on $\mathbb{P}^{n}$, say $\mathcal{F}\in\mathcal{L}(p_{1},...,p_{k};n)$. They are defined in homogeneous coordinates $z\in\mathbb{C}^{n+1}$ by a logarithmic 1-form of the type
@@
In this case, we have the following:
@@
In the statement, by generically transverse we mean that the set of tangencies between $V_{i}$ and $V_{j}$ has codimension $>2$.
As a consequence we have:
Corollary 2.5.
@@
1. In the case $tdg(\mathcal{G}_{\mathcal{F}})=1$ we take the foliation with first integral $F_{2}(x,y,z,t)=\frac{(xz+yt)y^{2}+x^{4}}{y^{4}}$, like in remark 2.7. As we have seen, this foliation has degree three and its Gauss map is birational.
@@
=$\mathcal{L}=\{\ell\in G_{1}(\mathbb{P}^{3})\,|\,\ell\cap Sing(\mathcal{F})=\emptyset\,\,\text{and $\ell$ has $d$ tangencies with $\mathcal{F}\}$}\,,$
where $G_{1}(\mathbb{P}^{3})$ denotes the grassmanian of lines of $\mathbb{P}^{3}$. Note that $\mathcal{L}$ is a Zariski open and dense subset of $G_{1}(\mathbb{P}^{3})$.
@@
1. A plane $P\in\mathcal{P}_{\ell}$, transverse to $Sing(\mathcal{F})$, has exactly one Morse tangency with $\mathcal{F}$, by proposition 3. This Morse tangency is in fact $\mathcal{G}_{\mathcal{F}}^{-1}(P)$, which we will denote as $\mathcal{M}_{\ell}(P)$.
@@
In particular, if we fix $\ell\in\mathcal{L}$ then the map $\mathcal{M}_{\ell}$ extends analytically to $\mathcal{P}_{\ell}$, so that the image $C_{\ell}:=\mathcal{M}_{\ell}(\mathcal{P}_{\ell})$ is a rational curve of ˘$\breve{\mathbb{P}}^{3}$.
@@
Claim 2.1.
The plane $P=\mathcal{G}_{\mathcal{F}}(p)$ is in the pencil $\mathcal{P}_{\ell}$. In particular, $\ell\subset P$. Moreover, $\mathcal{M}(P)=p$ and the curve $C_{\ell}$ is tangent to $P$ with tangent space $T_{p}C_{\ell}$ transverse to $T_{p}\ell$ (in the plane $P$).
@@
$f|_{z=\alpha y}(x,y)=\alpha y+x^{2}+y^{2}+a\alpha xy+b\alpha y^{2}+c\alpha^{2}y^{2}+h.o.t$
This implies that $\mathcal{M}(z=\alpha y)$ can be calculated as the zeroes of
@@
$=\left(2x+a\alpha y,\alpha+a\alpha x+(2+2b\alpha)y\right)+h.o.t$
@@
$\alpha\mapsto(X(\alpha),Y(a))=(o(\alpha^{2}),-\alpha/2+o(\alpha^{2}))\,.$
@@
$\mathcal{M}(z=\alpha y)=(o(\alpha^{2}),-\alpha/2+o(\alpha^{2}),-\alpha^{2}/2+o(\alpha^{3}))\,\implies$
$T_{p}C_{\ell}=(x=z=0)$ which is transverse to $T_{p}\ell$, in the plane $P$.∎Now, fix another point $q$ of tangency between $\ell$ and $\mathcal{F}$. Note that $\ell\subset\mathcal{G}_{\mathcal{F}}(q)\neq\mathcal{G}_{\mathcal{F}}(p)$. Since $T_{q}C_{\ell}$ is contained in $\mathcal{G}_{\mathcal{F}}(q)$ and transverse to $\ell$, we conclude that $T_{q}C_{\ell}$ is transverse to $P=\mathcal{G}_{\mathcal{F}}(p)$. This implies that $C_{\ell}$ cuts $P$ exactly in the $d$ points of tangency of $\ell$ with $\mathcal{F}$, but in just one of these points it is not transverse to $P$. Therefore, the degree of $C_{\ell}$ is $d+1$.
@@ 2.1. Problems
Here we state some problems that arise naturally from the results and examples.
@@
A good reference for the theory of k-webs is the book [32].
We have seen in remark 2.9 that the hypothesis of corollary 2.3 on foliations of $\mathbb{P}^{n}$ is not necessary for a Gauss map to be birational, although it is sufficient.
@@
We have seen that a foliation $\mathcal{F}$ of degree $d$ of $\mathbb{P}^{3}$ such that $tdg(\mathcal{G}_{\mathcal{F}})=d^{2}$ has a first integral. Then we can ask:
@@
In corollary 2.4 we have classified the generic logarithmic foliations on $\mathbb{P}^{3}$ whose Gauss map is birational.
@@
Is there such classification in the case of $\mathbb{P}^{n}$, $n\geq 4$.
We have seen in examples 9 and 11 that if $d=2$ or $d=3$ that $\mathcal{N}(d,3)=\{0,1,...,d^{2}\}$. A natural problem is the following:
@@ 3.1. Proof of theorem 1
A foliation of $\mathbb{P}^{2}$ of degree $d$ has at most $d^{2}$ non-degenerate singularities with trace zero. Moreover, if $\mathcal{F}\in\mathbb{F}ol(d;2)$ has $d^{2}$ non-degenerate singularities with trace zero then it has a first integral of the form $F/L^{d+1}$ where $deg(F)=d+1$ and $deg(L)=1$.
@@
1. $P(x,y)=p(x,y)+x\,g(x,y)$ and $Q(x,y)=q(x,y)+y\,g(x,y)$.
2. $max(deg(p),deg(q))=d$ and $g$ is homogeneous of degre $d$. Since $\ell_{\infty}$ is not $\mathcal{F}$-invariant we have $g\neq 0$.
@@
Proof. Let $V:=\overline{(P=0)}$ and $W:=\overline{(\Delta=0)}$ be the closures in $\mathbb{P}^{2}$ of $(P=0)$ and $(\Delta=0)$, respectively. Since $deg(P)=d+1$ and $deg(\Delta)=d$ the intersection $V\cap W$ contains $d^{2}+d$ points counted with multiplicity. Among these points we have the $d^{2}$ points $p_{1},...,p_{d^{2}}$. The other points of intersection are at the line of infinity $\ell_{\infty}$ and correspond to the $d$ linear factors of $g$. This implies the claim.∎
At this point, we can assert that $p_{1},...,p_{d^{2}}$ are all singularities of $\mathcal{F}$ with trace zero because $\Delta=0$ contains all such singularities. Therefore, the maximum number of non-degenerate singularities with zero trace is $d^{2}$: there is no other such singularity. We have proved above that $V$ and $W$ meet transversely in $\mathbb{C}^{2}$ at the points $p_{i}$, $1\leq i\leq d^{2}$, and at $\ell_{\infty}$ at the points corresponding to $g=0$. Since $Q(p_{i})=0$, $1\leq i\leq d^{2}$, and it meets $\ell_{\infty}$ and $W$ at the points corresponding to the $d$ linear factors of $g$, we obtain that $Q\in\left<P,\Delta\right>$, the ideal generated by $P$ and $\Delta$.
Therefore, we can write $Q=\alpha\,P+f\,\Delta$, where $\alpha\in\mathbb{C}^{*}$ and $deg(f)=1$. In particular, if $p\in Sing(\mathcal{F})\setminus\{p_{1},...,p_{d^{2}}\}$ then its trace is non zero and $f(p)=0$, because $P(p)=Q(p)=0$, but $\Delta(p)\neq 0$.
We have concluded that the $d+1$ singularities (counted with multiplicities) of $\mathcal{F}$ which are not in the set $\{p_{1},...,p_{d^{2}}\}$ must be in the line $(f=0)$. Since $\mathcal{F}$ has degree $d$ and $\overline{(f=0)}$ contains $d+1$ singularities of $\mathcal{F}$ the line $L:=\overline{(f=0)}$ must be $\mathcal{F}$-invariant (see [2]). Now, we change the affine coordinate system. Let $\mathbb{C}^{2}=\mathbb{P}^{2}\setminus L$. In this new affine coordinate system, the foliation can be represented by a form $\eta=A(x,y)dx+B(x,y)dy$, where $max(deg(A),deg(B))=d$, because the line at infinity $\ell_{\infty}=L$ is $\mathcal{F}$-invariant. In fact, since $\{p_{1},...,p_{d^{2}}\}\subset\mathbb{C}^{2}$ and $A(p_{i})=B(p_{i})=0$ for all $i=1,...,d^{2}$, we must have $deg(A)=deg(B)=d$.
Claim 3.2.
@@
Hence $\eta=dF$ where $deg(F)=d+1$. This implies that $\mathcal{F}$ has the first integral $F/L^{d+1}$. This finishes the proof of the claim and of proposition 5. ∎Let us finish the proof of theorem 1. Suppose that a foliation $\mathcal{F}$ of degree $d$ on $\mathbb{P}^{3}$ has a first integral of the form $F/L^{d+1}$, where $F\in\mathbb{C}[x,y,z,t]$ is homogeneous of degree $d+1$. If $F$ is non singular then for a generic hyperplane $\mathbb{P}^{2}\simeq\Sigma\subset\mathbb{P}^{3}$ the restriction $F|_{\Sigma}$ has $d^{2}$ Morse points, so that the fiber of the Gauss map associated has $d^{2}$ points. Conversely, let $\mathcal{F}$ be a foliation of $\mathbb{P}^{3}$ such that the generic fiber of the Gauss map has $d^{2}$ points. In particular, if $\Sigma\in\breve{\mathbb{P}}^{3}$ is a generic hyperplane then $\mathcal{F}|_{\Sigma}$ has $d^{2}$ singularties with trace zero, so that by proposition 5, $\mathcal{F}|_{\Sigma}$ has a first integral of the form $f/\ell^{d+1}$. This implies that $\mathcal{F}$ has a first integral of the form $F/L^{d+1}$ (see [26]). This finishes the proof of theorem 1. ∎

### v3

@@ 3. Since $deg(\mathcal{F}_{\omega})=n-1:=d$ we get
According to [10], the space of foliations of degree two on $\mathbb{P}^{n}$ where $n\geq 3$, has the following six irreducible components:
  • (a).
    PB(2;n)PB(2\,;n); the component containing pull-back of foliations of degree two on $\mathbb{P}^{2}$ by a linear map $L\colon\mathbb{P}^{n}\dasharrow\mathbb{P}^{2}$.
  • (b).
  • (c).
    $\mathcal{L}(1,1,1,1\,;n)$; component of foliations of degree two, defined by logarithmic forms of the type $\sum_{j=1}^{4}\lambda_{j}\frac{d\ell_{j}}{\ell_{j}}$, where $deg(\ell_{j})=1$, $1\leq j\leq 4$.
  • (d).
    $\mathcal{L}(1,1,2\,;n)$; component of foliations of degree two, defined by logarithmic forms of the type $\lambda_{1}\frac{d\ell_{1}}{\ell_{1}}+\lambda_{2}\frac{d\ell_{2}}{\ell_{2}}+\lambda_{3}\frac{dq}{q}$, where $deg(\ell_{1})=deg(\ell_{2})=1$ and $deg(q)=2$.
  • (e).
    $\mathcal{L}(1,3\,;n)$; component of foliations of degree two, with a first integral of the form $f/\ell^{3}$, where $deg(\ell)=1$ and $deg(f)=3$.
  • (f).
    $\mathcal{L}(2,2\,;n)$; component of foliations of degree two, with a first integral of the form $q_{1}/q_{2}$, where $deg(q_{1})=deg(q_{2})=2$.
Let $n\geq 4$, we will see that the Gauss maps of the generic foliations of degree two on $\mathbb{P}^{n}$ are not birational. Examples (a), (b), and (c), cannot be dominant on $\mathbb{P}^{n}$. For instance, an exceptional foliation on $\mathbb{P}^{n}$, is a linear pull-back of the exceptional foliation $\mathcal{E}$ on $\mathbb{P}^{3}$ (see [10]), and so its Gauss map cannot be dominant.
Generic foliations in components (d), (e), and (f), have dominant Gauss maps, but not birational. For instance, in case (e) we can assume that $f$ is a generic cubic polynomial. In homogeneous coordinates $(s,t,x,y,z)$ we can suppose that $\ell=s$, so that in the affine coordinate system $s=1$ the foliation $\mathcal{F}\in\mathcal{L}(1,3\,;n)$ has the first integral $f(1,t,x,y,z)$. The hyperplane $(t=c)$ is tangent to $\mathcal{F}$ at the points for which $f_{x}(1,c,x,y,z)=f_{y}(1,c,x,y,z)=f_{z}(1,c,x,y,z)=0$. If $f$ and $c$ are generic, this set contains $8$ points, whose image under the Gauss map is the hyperplane $(t=c)$, this implies that the map cannot be birational.
$f(s,t,x,y,z)=\frac{x^{2}+y^{2}+z^{2}}{st}:=\frac{q}{st}$
which is defined in homogeneous coordinates by the form
The Gauss map in the affine coordinate system $t=1$ is
and it has as inverse
$\mathcal{G}_{\mathcal{F}}^{-1}(a,b,c,e)=\left(-\frac{2a}{q(a,b,c)},-\frac{2b}{q(a,b,c)},-\frac{2c}{q(a,b,c)},\frac{1}{e},\right)$
then, the foliation $\mathcal{F}$ is self-dual.
Example 10.
In this example, we use the notation for the components of $\mathbb{F}ol(2;3)$, introduced in the previous one. Using Theorem 3 it is not difficult to see that there are two of these components for which the generic elements have birational Gauss map: the exceptional component of example 7 and the logarithmic component $\mathcal{L}(1,1,1,1;3)$ of example 3.
@@
This calculation can also be found in example 2.5 of [20]. We observe that $\mathcal{N}(2,3)=\{0,1,2,3,4\}$.
@@
Another example is a foliation in $\mathcal{L}(1,3;3)$ whose Gauss map is birational, is defined in homogeneous coordinates by the form $\omega=3\,f\,dz-z\,df$, where $f(x,y,z,t)=xz^{2}+yzt+y^{3}$.
In this case, we have $Sing(\mathcal{F}_{\omega})=(y=z=0):=V$ and $\mu(\mathcal{F}_{\omega},V)=6$, so that its Gauss map, $\mathcal{G}_{\omega}$, is birational according to Theorem 3. We have:
@@
The foliation associated with this map is $(\mathcal{G}_{\omega})_{*}(\mathcal{F}_{\omega})$ and it has the rational first integral:
@@
then, this is also a self-dual foliation.
Next, we will consider foliations defined by closed forms. Among them, we have logarithmic foliations on $\mathbb{P}^{n}$, say $\mathcal{F}\in\mathcal{L}(p_{1},...,p_{k};n)$. They are defined in homogeneous coordinates $z\in\mathbb{C}^{n+1}$ by a logarithmic 1-form of the type
@@
In this case, we have the following.
@@
In the statement, by generically transverse we mean that the set of tangencies between $V_{i}$ and $V_{j}$ has codimension >2>2.
As a consequence, we have:
Corollary 6.
@@
1. In the case $tdg(\mathcal{G}_{\mathcal{F}})=1$ we take the foliation with first integral $F_{2}(x,y,z,t)=\frac{(xz+yt)y^{2}+x^{4}}{y^{4}}$, like in remark 7. As we have seen, this foliation has degree three and its Gauss map is birational.
@@
[…]
where $G_{1}(\mathbb{P}^{3})$ denotes the Grassmannian of lines of $\mathbb{P}^{3}$. Note that $\mathcal{L}$ is a Zariski open and dense subset of $G_{1}(\mathbb{P}^{3})$.
@@
1. A plane $P\in\mathcal{P}_{\ell}$, transverse to $Sing(\mathcal{F})$, has exactly one Morse tangency with $\mathcal{F}$, because the Gauss map is birational. This Morse tangency is in fact $\mathcal{G}_{\mathcal{F}}^{-1}(P)$, which we will denote as $\mathcal{M}_{\ell}(P)$.
@@
In particular, if we fix $\ell\in\mathcal{L}$ then the map $\mathcal{M}_{\ell}$ extends analytically to $\mathcal{P}_{\ell}$, so that the image $C_{\ell}:=\mathcal{M}_{\ell}(\mathcal{P}_{\ell})$ is a rational curve of $\breve{\mathbb{P}}^{3}$.
@@
Claim 1.
The plane $P=\mathcal{G}_{\mathcal{F}}(p)$ is in the pencil $\mathcal{P}_{\ell}$. In particular, $\ell\subset P$. Moreover, $\mathcal{M}_{\ell}(P)=p$ and the curve $C_{\ell}$ is tangent to $P$ with tangent space $T_{p}C_{\ell}$ transverse to $T_{p}\ell$ (in the plane $P$).
@@
$f|_{z=\alpha y}(x,y,\alpha y)=\alpha y+x^{2}+y^{2}+a\alpha xy+b\alpha y^{2}+c\alpha^{2}y^{2}+h.o.t$
This implies that $\mathcal{M}_{\ell}(z=\alpha y)$ can be calculated as the zeroes of
@@
$\left(2x+a\alpha y,\alpha+a\alpha x+2(1+b\alpha+c\alpha^{2})y\right)+h.o.t$
@@
$\alpha\mapsto(X(\alpha),Y(\alpha))=\big(o(\alpha),-\frac{\alpha}{2}+o(\alpha)\big)\,.$
@@
$\mathcal{M}_{\ell}(z=\alpha y)=(o(\alpha),-\frac{\alpha}{2}+o(\alpha),-\frac{\alpha^{2}}{2}+o(\alpha^{2}))\,\implies$
$T_{p}C_{\ell}=(x=z=0)$ which is transverse to $T_{p}\ell$, in the plane PP.∎
Now, fix another point $q$ of tangency between $\ell$ and $\mathcal{F}$. Note that $\ell\subset\mathcal{G}_{\mathcal{F}}(q)\neq\mathcal{G}_{\mathcal{F}}(p)$. Since $T_{q}C_{\ell}$ is contained in $\mathcal{G}_{\mathcal{F}}(q)$ and transverse to $\ell$, we conclude that $T_{q}C_{\ell}$ is transverse to $P=\mathcal{G}_{\mathcal{F}}(p)$. This implies that $C_{\ell}$ cuts $P$ exactly in the $d$ points of tangency of $\ell$ with $\mathcal{F}$, but in just one of these points it is not transverse to $P$. Therefore, the degree of $C_{\ell}$ is $d+1$.
@@ 2.1. Problems
Here we state some problems that arise naturally from the results and examples discussed above.
@@
A good reference for the theory of k-webs is the book [33]. Here it is appropriate to include the following comment, which has been suggested by the referee: If $\mathcal{F}$ is an arbitrary codimension one foliation on $\mathbb{P}^{n}$ then $(\mathcal{G}_{\mathcal{F}})_{*}\mathcal{F}$ is the dual web or Legendre transform of $\mathcal{F}$ on $\breve{\mathbb{P}}^{n}$.
We have seen in remark 9 that the hypothesis of corollary 3 on foliations of $\mathbb{P}^{n}$ is not necessary for a Gauss map to be birational, although it is sufficient.
@@
We have seen in Theorem 1 that a foliation $\mathcal{F}$ of degree $d$ of $\mathbb{P}^{3}$ such that $tdg(\mathcal{G}_{\mathcal{F}})=d^{2}$ has a first integral. Then we can ask:
@@
In corollary 6 we have classified the generic logarithmic foliations on $\mathbb{P}^{3}$ whose Gauss map is birational.
@@
Is there such a classification in the case of $\mathbb{P}^{n}$, $n\geq 4$?
We have seen in examples 10 and $1$ that if $d=2$ or $d=3$ then $\mathcal{N}(d,3)=\{0,1,...,d^{2}\}$. A natural problem is the following.
@@ 3.1. Proof of Theorem 1
A foliation on $\mathbb{P}^{2}$ of degree $d$ has at most $d^{2}$ non-degenerate singularities with trace zero. Moreover, if $\mathcal{F}\in\mathbb{F}ol(d;2)$ has $d^{2}$ non-degenerate singularities with trace zero then it has a first integral of the form $F/L^{d+1}$ where $deg(F)=d+1$ and $deg(L)=1$.
@@
1. P(x,y)=p(x,y)+xg(x,y)P(x,y)=p(x,y)+x\,g(x,y) and Q(x,y)=q(x,y)+yg(x,y)Q(x,y)=q(x,y)+y\,g(x,y).
2. $max(deg(p),deg(q))=d$ and $g$ is homogeneous of degre $d$.
We can take the polynomials like this because the line $\ell_{\infty}=(z=0)$ is not $\mathcal{F}$-invariant, then $g\neq 0$ and $z\nmid g(x,y)$; therefore we can consider the homogeneous part of $g$ of degree d. The condition $max(deg(p),deg(q))=d$, comes from the fact that $\ell_{\infty}$ does not contain singularities of $\mathcal{F}$, then $z\nmid p(x,y)$ or $z\nmid q(x,y)$.
@@
Proof. Let $V:=\overline{(P=0)}$ and $W:=\overline{(\Delta=0)}$ be the closures in $\mathbb{P}^{2}$ of $(P=0)$ and $(\Delta=0)$, respectively. Since $deg(P)=d+1$ and $deg(\Delta)=d$ the intersection $V\cap W$ contains $d^{2}+d$ points counted with multiplicity. Among these points we have the $d^{2}$ points $p_{1},...,p_{d^{2}}$. The other points of intersection are at the line of infinity $\ell_{\infty}$ and correspond to the $d$ linear factors of $g$ (not necessarily different). This implies the claim.∎
At this point, we can assert that $p_{1},...,p_{d^{2}}$ are all singularities of $\mathcal{F}$ with trace zero because $\Delta=0$ contains all such singularities. Therefore, the maximum number of non-degenerate singularities with zero trace is $d^{2}$: there is no other such singularity.
We have proved above that $V$ and $W$ meet transversely in $\mathbb{C}^{2}$ at the points $p_{i}$, $1\leq i\leq d^{2}$, and at $\ell_{\infty}$ at the points corresponding to $g=0$. Since $Q(p_{i})=0$, $1\leq i\leq d^{2}$, and it meets $\ell_{\infty}$ and $W$ at the points corresponding to the $d$ linear factors of $g$, we obtain that $Q\in\left<P,\Delta\right$>, the ideal generated by $P$ and $\Delta$.
Therefore, we can write $Q=\alpha\,P+f\,\Delta$, where $\alpha\in\mathbb{C}^{*}$ and, since $deg(f)\leq 1$ in addition to the fact that $gcd(P,Q)=1$, we have that $deg(f)=1$. In particular, if $p\in Sing(\mathcal{F})\setminus\{p_{1},...,p_{d^{2}}\}$ then its trace is non zero and $f(p)=0$, because $P(p)=Q(p)=0$, but $\Delta(p)\neq 0$.
We have concluded that the $d+1$ singularities (counted with multiplicities) of $\mathcal{F}$, which are not in the set $\{p_{1},...,p_{d^{2}}\}$, must be in the line $(f=0)$. Since $\mathcal{F}$ has degree $d$ and $\overline{(f=0)}$ contains $d+1$ singularities of $\mathcal{F}$ the line $L:=\overline{(f=0)}$ must be $\mathcal{F}$-invariant (see [2]).
Now, we change the affine coordinate system. Let $\mathbb{C}^{2}=\mathbb{P}^{2}\setminus L$. In this new affine coordinate system, the foliation can be represented by a form $\eta=A(x,y)dx+B(x,y)dy$, where $max(deg(A),deg(B))=d$, because the line at infinity $\ell_{\infty}=L$ is $\mathcal{F}$-invariant. In fact, since $\{p_{1},...,p_{d^{2}}\}\subset\mathbb{C}^{2}$ and $A(p_{i})=B(p_{i})=0$ for all $i=1,...,d^{2}$, we must have $deg(A)=deg(B)=d$.
Claim 3.
@@
Hence $\eta=dF$ where $deg(F)=d+1$. This implies that $\mathcal{F}$ has the first integral $F/L^{d+1}$. This finishes the proof of the claim and of proposition 5. ∎
Let us finish the proof of Theorem 1. Suppose that a foliation $\mathcal{F}$ of degree $d$ on $\mathbb{P}^{3}$ has a first integral of the form $F/L^{d+1}$, where $F\in\mathbb{C}[x,y,z,t]$ is homogeneous of degree $d+1$. If $F$ is non singular then for a generic hyperplane $\mathbb{P}^{2}\simeq\Sigma\subset\mathbb{P}^{3}$ the restriction $F|_{\Sigma}$ has $d^{2}$ Morse points, so that the fiber of the Gauss map associated has $d^{2}$ points.
Conversely, let $\mathcal{F}$ be a foliation of $\mathbb{P}^{3}$ such that the generic fiber of the Gauss map has $d^{2}$ points. In particular, if $\Sigma\in\breve{\mathbb{P}}^{3}$ is a generic hyperplane then $\mathcal{F}|_{\Sigma}$ has $d^{2}$ singularties with trace zero, so that by proposition 5, $\mathcal{F}|_{\Sigma}$ has a first integral of the form $f/\ell^{d+1}$. This implies that $\mathcal{F}$ has a first integral of the form $F/L^{d+1}$ (see [12]). This finishes the proof of theorem 1. ∎